OpenStudy (anonymous):

Hi let me ask it another way, Solve the following equation, corrct to two decimal places: 4/x - 5 = 3x Any better?

OpenStudy (anonymous):

Hey I can do it ...May I

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

Hahah The best thing would be to multiply by x both sides \[4 - 5x = 3x^2\]

OpenStudy (anonymous):

Go on!

OpenStudy (anonymous):

Then some Manipulation \[3x^2 + 5x-4=0\] Now \[b^2 - 4ac > 0 => 25 + 48 \] So real roots \[x = \frac{ -5 \pm \sqrt{25 +48}}{6}\] Sorry for the delay I was unable to factor

OpenStudy (anonymous):

Not so easy:

OpenStudy (anonymous):

Now we need some Calculative skills\[x = 9.3 , .6\]

OpenStudy (anonymous):

Verify with Calculator

OpenStudy (anonymous):

Thank you